How many molecules are in 0.0021 mol of magnesium oxide?


1
Expert's answer
2021-04-07T11:48:00-0400

Solution:

magnesium oxide = MgO

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of MgO molecules = (0.0021 mol MgO)×(6.022×1023 molecules / 1 mol) = 4.78×1021 molecules

Number of MgO molecules = 1.26×1021 molecules


Answer: 1.26×1021 molecules are in 0.0021 mol of magnesium oxide (MgO).